OpenStudy (anonymous):

Using the given points, determine the slope. (0, 32) and (100, 212) slope = -9/5 slope = -5/9 slope = 5/9 slope = 9/5

OpenStudy (anonymous):

d

sam (.sam.):

\[\Huge Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

sam (.sam.):

\[\Huge Slope=\frac{212-32}{100-0}\] \[\Huge Slope=\frac{9}{5}\]